OpenStudy (anonymous):

Express 2a^-4 / 5b^-2 with positive exponents.

OpenStudy (welshfella):

Hint: a^-4 = 1 / a^4 and 1 / b^-2 = b^2

OpenStudy (anonymous):

\[\frac{2}{5 a^4 b^2} \]

OpenStudy (anonymous):

they gave me _/_ to put my answer

OpenStudy (welshfella):

2b^2/5a^4

OpenStudy (anonymous):

thank you

OpenStudy (welshfella):

yw